OpenStudy (anonymous):

If x and y are real numbers such that x > 1 and y < −1, then which of the following inequalities must be true? A. (x/y)> 1 B. abs(x^2)>abs(y) C. (x/3)− 5 >(y/3) − 5 D. x2 + 1 > y2 + 1 E. x^−2 > y^-2 ACT Practice says C, but I don't see why it can't be E

jimthompson5910 (jim_thompson5910):

If x = 10 and y = 3, then \[\Large x^{-2} > y^{-2}\] \[\Large 10^{-2} > 3^{-2}\] \[\Large \frac{1}{10^{2}} > \frac{1}{3^{2}}\] \[\Large \frac{1}{100} > \frac{1}{9}\] \[\Large 0.01 > 0.1111\] The last inequality is false, so the first inequality is false when x = 10 and y = 3. This counter example shows us that the original inequality is NOT true for all x and y such that x > 1 and y < -1.

OpenStudy (anonymous):

But y must be negative

OpenStudy (zarkon):

do what jim_thompson5910 did, but replace \(y=3\) with \(y=-3\)

OpenStudy (anonymous):

Then it would be true that 1/100>-1/9

jimthompson5910 (jim_thompson5910):

oh my bad, yeah y is negative. Zarkon has the proper fix

jimthompson5910 (jim_thompson5910):

\[\Large (-3)^{-2} = \frac{1}{(-3)^2} = \frac{1}{9}\]